OpenStudy (anonymous):

Having trouble factoring x^3 - 9x^2 + 27x - 27...I know the answer but don't know how to do it.

OpenStudy (anonymous):

I used the "difference of cubes"... not sure if there's a easier way... (x^3-27)-(9x^2-27x) =(x-3)(x^2+3x+9)-9x(x-3) =(x-3)(x^2+3x+9-9x) =(x-3)(x^2-6x+9) =(x-3)(x-3)(x-3) =(x-3)^3

OpenStudy (anonymous):

Thank you...that is very helpful. Wasn't thinking about re-grouping out of different order.